The rate at which drugs are metabolized depends on age. The elderly tend to metabolize drugs more slowly, whereas children metabolize some drugs more rapidly than adults. Diazepam is used to treat anxiety disorders and seizures in patients in all age groups. Its half-life in hours is estimated to be equal to the patient's age in years; in a 50 - year-old, diazepam would have approximately a 50-hour half-life. How long will it take for 95% of a dose of diazepam to be metabolized in a 5-year-old child compared to a 50-year-old adult, if we assume a first-order process? (be sure to use unit)
Answer #1

Dos LPJ tuz - mes of species & Syear old She R = 0.693 -0.693 – 0.13871.- Shr so year old te = 50ehr k = 0.693 - 1386 X10 soh